OpenStudy (anonymous):

How many solutions are there for the equation? x2 + 16 = 8x A. 2 B. 1 C. 0

OpenStudy (shayaan_mustafa):

2 solutions.

OpenStudy (anonymous):

how come ?

OpenStudy (shayaan_mustafa):

simple. look at the degree of eauation

OpenStudy (shayaan_mustafa):

it is 2 degree equation. 3 degree equations have 3 solutions. degree of equaions=no. of solutions for that equations.

Directrix (directrix):

x^2 - 8x + 16 = 0 (x-4)(x-4) = 0 x -4 = 0 or x - 4 = 0 x = 4 or x = 4 There is a double root but will appear as one on the graph. Based on the slant of the questions you've posted, I'm thinking that the question writer coded the answer to be 1. There is one distinct root.

Directrix (directrix):

Look back at this question. http://openstudy.com/users/alexisdelgado#/updates/4f71dc6be4b07738f5ae0329